OpenStudy (anonymous):

find the equation of the tangent line to the curve f(x) = (lnx/e^x) at x=1

OpenStudy (***[isuru]***):

differentiate the function u have.... can u do that ?

OpenStudy (anonymous):

im stuck on that part help

OpenStudy (anonymous):

do u use quotient rule

OpenStudy (***[isuru]***):

\[\frac{ d f(x) }{ dx } = \frac{ e^x \frac{ d \ln x }{ dx } - lnx \frac{ d e^x }{ dx }}{ (e^x)^{2} }\]

OpenStudy (***[isuru]***):

\[\frac{ d f(x) }{ dx } = \frac{ e^x \frac{ 1 }{ x } -( lnx \times e^x ) }{ (e^x)^{2} }\]

OpenStudy (anonymous):

f'(x)=\[f'\left( x \right)=\frac{ e ^{x}\frac{ 1 }{ x }-\ln x e ^{x} }{ e ^{2x} }\]

OpenStudy (anonymous):

ok thats what i got now what ??

OpenStudy (anonymous):

lost????

OpenStudy (***[isuru]***):

\[\frac{ d f(x) }{dx } = \frac{ e^x -(x \times lnx \times e^x ) }{ x (e^x)^{2} }\]

OpenStudy (***[isuru]***):

now substitute x= 1 in that equation

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

isnt it (e^x)^2 for the bottom ?

OpenStudy (***[isuru]***):

\[\frac{ d f(x) }{ dx} = \frac{ e^1 - ( 1 \times \ln 1 \times e^1) }{1 \times (e^1)^{2}}\]

OpenStudy (***[isuru]***):

i just make a common denominator by expanding e^x * (1/x) - ln x *e^x

OpenStudy (***[isuru]***):

e^1 = e ln 1 = 0

OpenStudy (anonymous):

when x=1, \[f'(1)=\frac{ e }{ e ^{2} }=\frac{ 1 }{ e }\] this is the slope at x=1 y=f(x)=ln1/e^1=0/e=0 point is (1,0)

OpenStudy (***[isuru]***):

which means \[\frac{ df(x) }{ dx } = \frac{ e }{ e^2 } = \frac{ 1 }{ e}\]

OpenStudy (***[isuru]***):

haha... u 've got help from 2 people at once... so i guess u have no doubts... XD

OpenStudy (anonymous):

yea i di thank you soo 1/e is the slope??/

OpenStudy (***[isuru]***):

u r welcome... and yes

OpenStudy (anonymous):

thank you isuri <3

OpenStudy (***[isuru]***):

hey.. im isuru... and isuri is a name for girls in our country :/

OpenStudy (anonymous):

awesoem where are you from ??

OpenStudy (***[isuru]***):

SL ...

OpenStudy (anonymous):

What lol

OpenStudy (***[isuru]***):

u know the asian country ...SRI LANKA

OpenStudy (anonymous):

never heard of it lol. I live in the U.S thats why ur teaching me lol

OpenStudy (***[isuru]***):

XD ... but it sad to hear that u have never herd of it .:(

OpenStudy (anonymous):

well i Google it lol. its small lol

OpenStudy (***[isuru]***):

it's a island... and a small one... but it's beautifull

OpenStudy (anonymous):

yea it should be lol. its like 10 degrees feranheit where im from

zepdrix (zepdrix):

Figure this one out alexis? c:

OpenStudy (anonymous):

good i am surjit and i am from punjab india living in USA

OpenStudy (anonymous):

actually the final answer asks me for the equation of a tangent line. and my answer and multiple choice don't match ):

zepdrix (zepdrix):

uh ohs! :U

OpenStudy (***[isuru]***):

wt a the choices ?

zepdrix (zepdrix):

So you found the `slope` of your tangent line. Are the choices given in point-slope form `y-yo = m(x-xo)` or in slope-intercept form `y=mx+b`

OpenStudy (anonymous):

y-0=1/e(x-1) y=1/e(x-1)

OpenStudy (anonymous):

(a) x-ey-1=0 b x+ey-y=0 c x-y-1=0 d ex+y-1=0 e ex-y-1=0

OpenStudy (anonymous):

ey=x-1 orx-ey-1=0

OpenStudy (anonymous):

ok suri than you <3 im done bothering people lol

OpenStudy (anonymous):

a is the answer.

OpenStudy (anonymous):

thanks your the best

OpenStudy (anonymous):

yw